OpenStudy (anonymous):

Please help me Solve for x: 4|2x - 2| + 10 = 34. x = -4, x = 2 x = 4, x = -4 x = 2, x = -2 x = 4, x = -2

OpenStudy (anonymous):

@sammixboo

OpenStudy (jdoe0001):

\(\large { 4|2x - 2| + 10 = 34\implies 4|2x - 2|= 34-10 \\ \quad \\ |2x - 2|= \cfrac{34-10}{4}\to \begin{cases} +(2x - 2)= \cfrac{34-10}{4} \\ \quad \\ -(2x - 2)= \cfrac{34-10}{4} \end{cases} }\) solve each of the two scenarios, for "x"

OpenStudy (anonymous):

i got it! :D can you help me out with one more @jdoe0001